# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v3

@@ B.4 Modified log-linear learning in symmetric potential games
where the suboptimality gap $\Delta$ is given by $\Delta:=\min_{x\in\Psi^{\mathcal{A}}:\Phi(x)<\Phi(x^{*})}\left(\Phi(x^{*})-\Phi(x)\right)$ with $x^{*}\in\Psi_{*}^{\mathcal{A}}$. Then, for
$\beta\geq$ $\frac{1}{\max\{\epsilon,\Delta\}}\log\bigg{(}(N+1)^{A-1}\bigg{(}\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\bigg{)}\bigg{)},$ (23)
it holds that
$\frac{Y_{*}}{Y(\epsilon)+(Y-Y(\epsilon))e^{-\beta\min\{\epsilon,\Delta\}}}\geq 1-\epsilon,$
@@
where we used that $Y\leq(N+1)^{A-1}$. We deduce that for $\beta=\Omega(\frac{1}{\max\{\epsilon,\Delta\}}\log(\frac{N^{A}}{\epsilon}))$, it holds that:
$\mathbb{E}_{x\sim\mu_{m}}[\Phi(x)]\geq(1-\epsilon)\max_{x\in\Psi^{\mathcal{A}}}\Phi(x)\geq\max_{x\in\Psi^{\mathcal{A}}}\Phi(x)-\epsilon.$
@@
The proof now follows from the same analysis as in the proof of Theorem $3$ in (Shah and Shin, 2010) with the exception that we do not make use of Lemma $6$ in (Shah and Shin, 2010) but replace it with our analysis above. Concretely, we set $\beta$ as specified in Equation (23) rather than as in (Shah and Shin, 2010, Eq. (8)).
@@ Appendix C Binary log-linear learning
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The additional term $1/A$ stems from the fact that player $i$ first randomly samples an action $\tilde{a}_{i}$ and then decides between this action and her previous action. Arslan et al. (2007) show that its stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ is given by :
@@
The stationary distribution of binary log-linear learning is the same as that of log-linear learning (Equation (5)). Thus, log-linear- and binary log-linear learning converge to an approximately efficient Nash equilibrium in the long run. We briefly outline the proof of Theorem 4.1 and then provide a detailed proof.
Proof C.1.
The proof follows from the same line of arguments as the proof of Theorem 3.1. In particular, the first step in the proof of Theorem 3.1 remains the same since binary log-linear learning has the same stationary distribution as log-linear learning. In the second step in the proof of Theorem 3.1, the main difference is that the transition matrix in (24) of binary log-linear learning differs from the transition matrix in (4) of log-linear learning. Thus, the log-Sobolev constant of binary log-linear can be lower-bounded as follows:
@@
Then, we use Lemma 2.2 to show that
@@
Proof C.2.
of Theorem 4.1 By Lemma 3.4, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
@@
where we used that by definition of $P$ in (24) and $\mu$ in (25) $\mu_{\min}$ and $p_{\min}$ can be lower-bounded as follows:
@@
Equation (2) in Lemma 2.2 provides the following upper bound on the mixing time:
@@
$\beta$ $\geq\frac{1}{\max\{\epsilon/2,\Delta\}}\log\bigg{(}(A^{N}-A^{N}(\epsilon/2))\bigg{(}\frac{4}{\epsilon A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\bigg{)}\bigg{)}.$ (28)
@@ Appendix D Robustness of log-linear learning
In this section, we prove Theorem 4.3 and apply it to the corrupted-utility case to prove Corollary 4.4.
D.1 Proof of Theorem 4.3
Proof D.1.
Consider a learning rule $P$, to prove Theorem 4.3, we first provide a decomposition that relates the expected value of the potential when the agents follow $P$ defined in Equation (8) to the same quantity when the agents follow $P_{\ell}$ defined in Equation (4) instead.
We have for all $t,t^{\prime}\in\mathbb{N}$ that:
$\mathbb{E}_{\mu_{0}P^{t}}[\Phi]$ $=\mathbb{E}_{\mu_{0}P_{\ell}^{t^{\prime}}}[\Phi]+\mathbb{E}_{\mu_{0}P^{t}}[\Phi]-\mathbb{E}_{\mu_{0}P_{\ell}^{t^{\prime}}}[\Phi]$
$\geq\mathbb{E}_{\mu_{0}P_{\ell}^{t^{\prime}}}[\Phi]-\sqrt{A^{N}}\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ (29)
where the last line follows because $|\Phi(a)|\leq 1$ for all $a\in\mathcal{A}^{N}$ and because .$\|.\|_{1}\leq\sqrt{A^{N}}\|\cdot\|_{2}$. The rest of the proof is based on controlling $\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ using our mixing time results.
Decomposition: Using Lemma 4.2 we obtain:
$\|P^{t}-P_{\ell}^{t^{\prime}}\|_{2}$ $\leq\|P^{t}-\mu\|_{2}+\|P_{\ell}^{t^{\prime}}-\mu_{\ell}\|_{2}+\|\mu-\mu_{\ell}\|_{2}$
$\leq 2\|P^{t}-\mu\|_{TV}+\|P_{\ell}^{t^{\prime}}-\mu_{\ell}\|_{2}+L(P_{\ell})\|P-P_{\ell}\|_{2},$
where $L(P_{\ell})=\frac{2A^{N}}{\rho(P_{\ell}P_{\ell}^{*})}(\log\log\frac{1}{\mu_{\ell,\min}}+\log(8A^{N}))$ follows from Lemma 4.2. In Theorem 3.1, we showed that $\mu_{\ell,\min}\geq\frac{e^{-\beta}}{A^{N}}$ and $\rho(P_{\ell}P_{\ell}^{*})\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}}$, therefore
$L(P_{\ell})\leq\frac{25N^{2}A^{N+5}e^{4\beta}}{8\pi^{2}}(\log\log A^{N}e^{\beta}+\log(8A^{N})).$
Thus, for
$\begin{cases}t&\geq t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}})))\\ t^{\prime}&\to\infty\\ \beta&=\log\bigg{(}(A^{N}-A^{N}(\epsilon/2))\bigg{(}\frac{4}{\epsilon/2A^{N}(\epsilon/2)}-\frac{1}{A^{N}(\epsilon/2)}\bigg{)}\bigg{)}/\max\{\epsilon/2,\Delta\}\end{cases}$
we have
$\begin{cases}\|P^{t}-P_{\ell}^{t}\|_{2}&\leq\epsilon/\left(2\sqrt{A^{N}}\right)+L(P_{\ell})\|P-P_{\ell}\|_{2}\\ \mathbb{E}_{a\sim\mu^{0}P_{\ell}^{t^{\prime}}}[\Phi(a)]&\geq\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon/2,\\ L(P_{\ell})&=\mathcal{O}\Bigg{(}N^{2}A^{N+5}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}\Bigg{(}\log\log A^{N}e^{\frac{\log(A^{N}/\epsilon)}{\max\{\epsilon,\Delta\}}}+\log(A^{N})\Bigg{)}\Bigg{)},\end{cases}$
where the second line follows from Lemma 3.3. Plugging the above inequalities with the bound on $L$ from Lemma 4.2 into the decomposition (29) proves the desired result. We now bound the mixing time $t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}}))$ of the Markov chain induced by $P$.
Mixing time bound: To bound the mixing time of the Markov chain induced by $P$, we use inequality (2) and Lemma 3.4. Assuming a lower bound of $p_{\min}/N$ on the probabilities of all feasible transitions implies a bound on the stationary distribution as we will show next.
Lower bound $(\mu_{P})_{\min}$: Since $P$ has a positive probability of transitioning from $a\in\mathcal{A}^{N}$ to any $\tilde{a}\in\mathcal{N}(a)$, it follows that the corresponding $N$-step transition $P^{N}$ has a positive probability of transitioning from any $a\in\mathcal{A}^{N}$ to any $a^{\prime}\in\mathcal{A}^{N}$, i.e.,
Note that the least probable transitions are when $a$ and $a^{\prime}$ are such that $\forall i\in[N]:\>a_{i}\neq a^{\prime}_{i}$. For all such transitions, the possible paths using $P^{N}$ are the permutations of $\{1,\ldots,N\}$ (each of the $N$ steps is a new player updating their action). There are $N!$ such permutations and each player $i\in[N]$ can update $a_{i}$ to $a^{\prime}_{i}$ with probability larger than $p_{\min}/N$.
Since $P$ is an irreducible and aperiodic transition matrix, the Markov chain induced by $P$ has a unique stationary distribution $\mu_{P}$. It is known that the Markov chain induced by $P^{N}$ has the same stationary distribution as the one induced by $P$. Therefore, we have for all $a\in\mathcal{A}^{N}$:
$\mu_{P}(a)$ $=\sum_{\tilde{a}\in\mathcal{A}^{N}}P^{N}_{\tilde{a},a}\mu_{P}(\tilde{a})$
$\geq\sum_{\tilde{a}\in\mathcal{A}^{N}}N!\>(p_{\min}/N)^{N}\mu_{P}(\tilde{a})=N!\>(p_{\min}/N)^{N}.$
Therefore, $(\mu_{P})_{\min}\geq N!\>(p_{\min}/N)^{N}$.
Deducing the mixing-time bound: We can now give an explicit bound on the mixing time of the transition $P$. First, we have by Lemma 3.4:
$\rho(PP^{*})$ $\geq\frac{16\pi^{2}A^{N}(\mu_{P})_{\min}p_{\min}^{3}}{25N^{2}}$
$\geq\frac{4\pi^{2}A^{N}p_{\min}^{N+3}N!}{25N^{N+2}}.$
Then, using Stirling’s formula, we have $N!\geq\sqrt{2\pi N}\left(\frac{N}{e}\right)^{N}$, therefore
$\rho(PP^{*})\geq\frac{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}{25N^{3/2}e^{N}}$
Then, using inequality (2), we have:
$t_{\operatorname{mix}}^{P}(\epsilon/(4\sqrt{A^{N}}))$ $\leq\frac{1}{\rho(PP^{*})}\Bigg{(}\log\log\frac{1}{(\mu_{P})_{\min}}+2\log\frac{4\sqrt{A^{N}}}{\epsilon}\Bigg{)}$
$\leq\frac{25N^{3/2}e^{N}}{(2\pi)^{5/2}A^{N}p_{\min}^{N+3}}\Bigg{(}\log\log\frac{e^{N}}{p_{\min}^{N}\sqrt{2\pi N}}+2\log\frac{4\sqrt{A^{N}}}{\epsilon}\Bigg{)}.$
This concludes the proof of Theorem 4.3.
@@
Proof D.2.
If all players adhere to log-linear learning with corrupted utilities, the induced Markov chain’s transition matrix $\hat{P}$ is given, for all $a,\tilde{a}\in\mathcal{A}^{N}$ by:
@@
Finally, since $2\beta\xi\leq 1$ and by using that: $01-e^{-x}<x\text{ for }x>$ 0, and that: $e^{x}-1<\frac{7}{4}x\text{ for }x\in[0,1]$. Then,
@@
$\|\hat{P}-P\|_{2}$ $\leq\sqrt{\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\frac{49}{4N^{2}}\beta^{2}\xi^{2}}$
$=\frac{7A^{N}}{2N}\beta\xi.$
@@
Proof D.3.
If all players adhere to fixed-share log-linear learning, the induced Markov chain’s transition matrix $\hat{P}$ is given, for all $a,\tilde{a}\in\mathcal{A}^{N}$ by:
@@
$\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\right)^{2}$ $=\frac{\xi^{2}}{N^{2}}\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\frac{1}{A^{2}}-\frac{2e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{A\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}+\frac{e^{2\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\left(\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}\right)^{2}}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$
@@
[…]
@@
$\|\hat{P}-P\|_{2}$ $\leq\sqrt{\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\right)^{2}}$
$\leq\xi\sqrt{\frac{A^{N}}{N}}.$
@@
D.4 Proof of Lemma 4.2
Proof D.4.
Denote by $M\in\mathbb{R}^{A^{N}\times A^{N}}$ the matrix, where each row corresponds to $\mu_{1}$. For all $t\in\mathbb{N}$, we have that:
$\mu_{1}-\mu_{2}$ $=(P_{1}^{t})^{\top}(\mu_{1}-\mu_{2})+(P_{1}^{t}-P_{2}^{t})^{\top}\mu_{2}$
$=(P_{1}^{t}-M)^{\top}(\mu_{1}-\mu_{2})+M^{\top}(\mu_{1}-\mu_{2})$
$\quad+(P_{1}^{t}-P_{2}^{t})^{\top}\mu_{2}.$
This yields:
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq\|(P_{1}^{t}-M)^{\top}(\mu_{1}-\mu_{2})\|_{2}$
$\quad+\|M^{\top}(\mu_{1}-\mu_{2})\|_{2}+\|(P_{1}^{t}-P_{2}^{t})^{\top}\|_{2}\|\mu_{2}\|_{2}$
$\leq\|P_{1}^{t}-M\|_{2}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|M^{\top}(\mu_{1}-\mu_{2})\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}$
$\leq 2\sqrt{A^{N}}\|P_{1}^{t}-M\|_{TV}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|M^{\top}(\mu_{1}-\mu_{2})\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}$
where in the last inequality we used the equivalence of $\|\cdot\|_{2}$ and $\|\cdot\|_{1}$ and that $\|\cdot\|_{1}=2\|\cdot\|_{TV}$ by definition of the total variation distance. Furthermore:
$M^{\top}(\mu_{1}-\mu_{2})$ $=\bigg{(}\mu_{1}(a)\underbrace{\sum_{a^{\prime}\in\mathcal{A}^{N}}(\mu_{1}(a^{\prime})-\mu_{2}(a^{\prime}))}_{=0}\bigg{)}_{a\in\mathcal{A}^{N}}$
$=0.$
Therefore, we obtain that:
$\quad+\|P_{1}^{t}-P_{2}^{t}\|_{2}.$ (31)
Note that for the second term on the right-hand side in the equation above we have:
$=\sum_{l=1}^{t}(P_{1}^{t-l+1}P_{2}^{l-1}-P_{1}^{t-l}P_{2}^{l})$
$=\sum_{l=1}^{t}(P_{1}^{t-l}(P_{1}-P_{2})P_{2}^{l-1}).$
By applying the norm operator we find that:
$\leq tA^{N}\|P_{1}-P_{2}\|_{2},$
since $\|P\|_{2}\leq\sqrt{A^{N}}$ holds for all transition matrices $P$ over $\mathcal{A}^{N}$, and in particular for $P_{1}^{t-l}$ and $P_{2}^{l-1}$. By plugging the above into inequality (31) we find:
$\quad+tA^{N}\|P_{1}-P_{2}\|_{2}.$
The proof is then concluded by using the mixing-time bound from inequality (2).
Appendix E Numerical Illustration
We illustrate our convergence time results for log-linear learning for two-player identical interest games. In our setup, each player has the action set $\mathcal{A}=\{1,2,\ldots,10\}$, and the players have the same utility matrix denoted by $\{U(a_{1},a_{2})\}_{a_{1},a_{2}\in[10]}$. We simulate $10$ different utility matrices $U(\cdot,\cdot)\in[0,1]^{10\times 10}$ of the following form: we first fix a suboptimality gap $\Delta$, then we set $U(1,1)=1$, $U(10,10)=1-\Delta$, and the remaining entries of the utility matrix are sampled uniformly from the range $[0,1-\Delta)$.
Figure 1: Expected potential value when all players follow log-linear learning with $\beta$ set as in Equation (6). Left: performance for $\epsilon=0.05$ and different values of $\Delta$. Right: performance for $\Delta=0.1$ and different values $\epsilon$. Lines are averages over 10 randomly generated games, shaded areas represent one standard deviation, and the stars mark the first time the desired precision $1-\epsilon$ is reached.
In our first experiment, we fix $\epsilon$ as $0.05$ and vary the suboptimality gap $\Delta$ in $[0.15,0.10,0.075]$. We plot the expected potential value as a function of time when all players follow log-linear learning with temperature parameter $\beta$ set as in Equation (6). Figure 1 (left) demonstrates that the convergence time increases as the suboptimality gap $\Delta$ decreases. This shows that the suboptimality gap $\Delta$ quantifies the convergence time of log-linear learning well and therefore should be taken into account when bounding the convergence time, as done in our Theorem 3.1. In addition, we observe that this increase in convergence time is super-linear which aligns with the polynomial convergence time of $\tilde{\mathcal{O}}((\frac{1}{\epsilon})^{1/\Delta})$ that we proved in Theorem 3.1 when $\epsilon$ is smaller than $\Delta$.
In our second experiment, we fix $\Delta$ as $0.1$ and vary the precision level $\epsilon$ in $[0.1,0.05,0.025,0.02]$. We then plot the expected potential value as a function of time when all players follow log-linear learning with temperature parameter $\beta$ set as in Equation (6). Figure 1 (right) shows that the convergence time increases as the value of $\epsilon$ decreases. In other words, the convergence time of log-linear learning depends inversely on the desired suboptimality. This is also reflected in our convergence time bounds of order $\tilde{\mathcal{O}}((\frac{1}{\epsilon})^{1/\max\{\epsilon,\Delta\}})$ in Theorem 3.1.

### v4

@@ Appendix B Convergence of log-linear learning
where the suboptimality gap $\Delta$ is given by $\Delta:=\min_{a\in\mathcal{A}^{N}:\Phi(a)<\Phi(a^{*})}\left(\Phi(a^{*})-\Phi(a)\right)$ with $a^{*}\in\mathcal{A}^{N}_{*}$. If
$\beta\geq\frac{1}{\Delta}\log((A^{N}-A^{N}_{*})(\frac{1}{\epsilon A_{*}^{N}}-\frac{1}{A_{*}^{N}})),$ (28)
then $\frac{A^{N}_{*}}{A^{N}_{*}+(A^{N}-A^{N}_{*})e^{-\beta\Delta}}\geq 1-\epsilon$. If $\beta$ is set as in Equation (28), injecting the last inequality into Equation (27) implies:
$\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq(1-\epsilon)\Phi(a^{*})=\Phi(a^{*})-\epsilon,$
where we used that $\Phi(a^{*})\leq 1$. This concludes the proof.
B.2 Proof of Corollary 3.3
Define the set $\Psi_{*}^{\mathcal{A}}=\{x^{*}\in\Psi^{\mathcal{A}}\>|\>x^{*}\in\operatorname*{arg\,max}_{x\in\Psi^{\mathcal{A}}}\!\!\Phi_{m}(x)\}$ as the set of potential maximizers with cardinality $Y_{*}\!=\!|\Psi_{*}^{\mathcal{A}}|$. Then, the expected value of the potential function $\Phi_{m}$ over the stationary distribution $\mu_{m}$ of modified log-linear learning in (4) can be bounded as follows:
$\mathbb{E}_{x\sim\mu}[\Phi_{m}(x)]=\sum_{x\in\Psi^{\mathcal{A}}}\frac{e^{\beta\Phi_{m}}(x)}{\sum_{\tilde{x}\in\Psi^{\mathcal{A}}}e^{\beta\Phi_{m}}(\tilde{x})}\Phi_{m}(x)$ (29)
$\geq\!\!\!\!\sum_{x\in\Psi_{*}^{\mathcal{A}}}\!\!\frac{\Phi_{m}}{(}x){\sum_{\tilde{x}\in\Psi_{*}^{\mathcal{A}}}e^{\beta(\Phi_{m}(\tilde{x})-\Phi_{m}(x))}\!+\!\sum_{\tilde{x}\in\Psi^{\mathcal{A}}\!\setminus\!\Psi_{*}^{\mathcal{A}}}e^{\!\beta(\Phi_{m}(\tilde{x})-\Phi_{m}(x))}}$
$\geq\sum_{a\in\Psi_{*}^{\mathcal{A}}}\frac{\Phi_{m}(x)}{\sum_{\tilde{x}\in\Psi_{*}^{\mathcal{A}}}e^{0}+\sum_{\tilde{x}\in\Psi^{\mathcal{A}}\setminus\Psi_{*}^{\mathcal{A}}}e^{-\beta\Delta}}$
$\geq\frac{Y_{*}}{Y_{*}+(Y-Y_{*})e^{-\beta\Delta}}\Phi_{m}(x^{*}),$
where $\Delta:=\min_{x\in\Psi^{\mathcal{A}}:\Phi_{m}(x)<\Phi_{m}}(x^{*})\left(\Phi_{m}(x^{*})-\Phi_{m}(x)\right)$ is the suboptimality gap with $x^{*}\in\Psi_{*}^{\mathcal{A}}$. Then, for
$\beta\geq$ $\frac{1}{\Delta}\log\bigg((N+1)^{A-1}\bigg(\frac{1}{\epsilon Y_{*}}-\frac{1}{Y_{*}}\bigg)\bigg),$ (30)
It holds that:
$\frac{Y_{*}}{Y_{*}+(Y-Y_{*})e^{-\beta\Delta}}\geq 1-\epsilon,$
@@
where we used that $Y\leq(N+1)^{A-1}$. We deduce that for $\beta=\Omega\left(\frac{1}{\Delta}\log\left(\frac{N^{A}}{\epsilon}\right)\right)$, it holds that:
$\mathbb{E}_{x\sim\mu_{m}}[\Phi_{m}(x)]\geq(1-\epsilon)\max_{x\in\Psi^{\mathcal{A}}}\Phi_{m}(x)\geq\max_{x\in\Psi^{\mathcal{A}}}\Phi_{m}(x)-\epsilon.$
@@
The proof now follows from the same analysis as in the proof of Theorem $3$ in [38] with the exception that we replace Lemma 66 in [38] with our analysis above. Concretely, we set $\beta$ as specified in Equation (30) rather than as in [38, Eq. (8)].
@@ Appendix C Binary log-linear learning
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. The additional term $1/A$ stems from the fact that player $i$ first randomly samples an action $\tilde{a}_{i}$ and then decides between this action and her previous action. [1] show that its stationary distribution $\mu\in\Delta(\mathcal{A}^{N})$ is given by:
@@
Importantly, note that the stationary distribution of binary log-linear learning is the same as that of log-linear learning (Equation (4)). Thus, log-linear and binary log-linear learning converge to an approximately efficient Nash equilibrium in the long run. We briefly outline the proof of Theorem 5.1 and then provide a detailed proof.
Proof outline:
The proof follows from the same line of arguments as in the proof of Theorem 3.1. In particular, the first step in the proof of Theorem 3.1 remains the same since binary log-linear learning has the same stationary distribution as log-linear learning. Compared to the second step in the proof of Theorem 3.1, the main difference is that the transition matrix defined in (31) of binary log-linear learning differs from the transition matrix defined in (3) of log-linear learning. Thus, the log-Sobolev constant of binary log-linear can be lower-bounded as follows:
@@
Then, we use Lemma A.1 to show that
@@
Proof (Theorem 5.1) By Lemma 4.2, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
@@
where we used that by definition of $P$ in Equation (31) and $\mu$ in Equation (32), $\mu_{\min}$ and $p_{\min}$ can be lower-bounded as follows:
@@
Equation (23) in Lemma A.1 provides the following upper bound on the mixing time:
@@
$\beta\geq\frac{1}{\Delta}\log\left(\left(A^{N}-A_{*}^{N}\right)\left(\frac{1}{\epsilon A_{*}^{N}}-\frac{1}{A_{*}^{N}}\right)\right)$
@@ Appendix D Robustness of log-linear learning
In this section, we provide a prood of Lemma 5.3, Corollaey 5.4, and Corollary 5.5.
D.1 Proof of Lemma 5.3
Denote by $M\in\mathbb{R}^{A^{N}\times A^{N}}$ a matrix where each row corresponds to $\mu_{1}$. For all $t\in\mathbb{N}$, we have that:
$\mu_{1}-\mu_{2}$ $=\langle P_{1}^{t},\mu_{1}-\mu_{2}\rangle+\langle P_{1}^{t}-P_{2}^{t},\mu_{2}\rangle$
$=\langle P_{1}^{t}-M,\mu_{1}-\mu_{2}\rangle+\langle M,\mu_{1}-\mu_{2}\rangle+\langle P_{1}^{t}-P_{2}^{t}),\mu_{2}\rangle.$
This yields:
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq\|\langle P_{1}^{t}-M,\mu_{1}-\mu_{2}\rangle\|_{2}$
$\quad+\|\langle M,\mu_{1}-\mu_{2}\rangle\|_{2}+\|(P_{1}^{t}-P_{2}^{t})^{\top}\|_{2}\|\mu_{2}\|_{2},$
then,
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq\|P_{1}^{t}-M\|_{2}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|\langle M,\mu_{1}-\mu_{2}\rangle\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}$
$\leq 2\sqrt{A^{N}}\|P_{1}^{t}-M\|_{TV}\|\mu_{1}-\mu_{2}\|_{2}$
$\quad+\|\langle M,\mu_{1}-\mu_{2}\rangle\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}$
where in the last inequality we used the equivalence of $\|\cdot\|_{2}$ and $\|\cdot\|_{1}$ and that by definition of the total variation distance $\|\cdot\|_{1}=2\|\cdot\|_{TV}$. Furthermore:
$\langle M,\mu_{1}-\mu_{2}\rangle=\big(\mu_{1}(a)\sum_{a^{\prime}\in\mathcal{A}^{N}}(\mu_{1}(a^{\prime})-\mu_{2}(a^{\prime}))\big)_{a\in\mathcal{A}^{N}}=0.$
Therefore, we obtain that:
$\|\mu_{1}-\mu_{2}\|_{2}$ $\leq 2\sqrt{A^{N}}\|P_{1}^{t}-M\|_{TV}\|\mu_{1}-\mu_{2}\|_{2}+\|P_{1}^{t}-P_{2}^{t}\|_{2}.$ (35)
For the second term in the equation above, we have:
$=\sum_{l=1}^{t}(P_{1}^{t-l}(P_{1}-P_{2})P_{2}^{l-1}).$
By applying the norm operator and since $\|P\|_{2}\leq\sqrt{A^{N}}$ holds for all $P$ over $\mathcal{A}^{N}$ including $P_{1}^{t-l}$ and $P_{2}^{l-1}$ we find that:
$\leq tA^{N}\|P_{1}-P_{2}\|_{2}.$
Plugging the above in Inequality (35) we obtain:
$\quad+tA^{N}\|P_{1}-P_{2}\|_{2}.$
We conclude using the mixing-time bound of Inequality (23).
@@
Proof If all players adhere to log-linear learning with corrupted utilities, the induced Markov chain’s transition matrix $\hat{P}$ is given, for all $a,\tilde{a}\in\mathcal{A}^{N}$ by:
@@
Finally, since $2\beta\xi\leq 1$ and by using that: $1-e^{-x}<x\text{ for }x>0$, and that: $e^{x}-1<\frac{7}{4}x\text{ for }x\in[0,1]$. Then,
@@
$\|\hat{P}-P\|_{2}$ $\leq\sqrt{\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\frac{49}{4N^{2}}\beta^{2}\xi^{2}}=\frac{7A^{N}}{2N}\beta\xi.$
@@
Proof If all players adhere to fixed-share log-linear learning, the induced Markov chain’s transition matrix $\hat{P}$ is given, for all $a,\tilde{a}\in\mathcal{A}^{N}$ by:
@@
$\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\right)^{2}$
$=\frac{\xi^{2}}{N^{2}}\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\frac{1}{A^{2}}-\frac{2e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{A\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}+\frac{e^{2\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\left(\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}\right)^{2}}\right)\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$
@@
where the second line follows because from any action profile $a\in A^{N}$, there are $NA$ possible transitions ($A$ possible actions times $N$ possible player selections). We also used $\sum_{\tilde{a}\in\mathcal{A}^{N}}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}=1$ and that $\sum_{\tilde{a}\in\mathcal{A}^{N}}\frac{e^{\beta U_{i}(\tilde{a}_{i},\tilde{a}_{-i})}}{\left(\sum_{a_{i}^{\prime}\in\mathcal{A}}e^{\beta U_{i}(a_{i}^{\prime},\tilde{a}_{-i})}\right)^{2}}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}\leq 1$.
@@
$\|\hat{P}-P\|_{2}$ $\leq\sqrt{\sum_{a,\tilde{a}\in\mathcal{A}^{N}}\left(\hat{P}_{a,\tilde{a}}-P_{a,\tilde{a}}\right)^{2}}\leq\xi\sqrt{\frac{A^{N}}{N}}.$